# Graded lithium-ion battery pouch cells to homogenise current distributions and mitigate lithium plating

### v1

@@ 2.6 Model validation
Figure 2(b) shows that the model was able to capture the evolution of the maximum, minimum, and average values of the cell temperature across the plane of the pouch cell during the cycling. A main point of difference between the trajectories of the model and the experimental data is that the model’s trajectories are much smoother. The increased stochasticity of the data is to be expected, as the sensing noise and the inherent randomness in the microstructure introduced by manufacturing limitations are not captured by Eqn. (3). However, the model is still able to capture the general trends of the data, with the time constants and steady-state values of the first-order responses seen in the temperature dynamics being captured.
Figure 3 compares snapshots in the plane of the temperature distribution of the model and the experimental data 20, where the snapshots were taken at $t=$ 100s, 500s, 1000s, and 2500s. Even though the underlying model of Eqn. (3) is simple, strong agreement was made with the experimental thermal data from Lin, et al, 202220. In particular, the formation of a thermal hot spot as well as the general shape of the temperature contours were captured by the model, although, again, the model’s diffusion dynamics resulted in smoother contours than those seen in the data.
Figure 4 illustrates how the validated model could be used to infer the impact of design changes in the pouch cell on the degree of spatial heterogeneity across the plane. Defining the variation in the current density as the difference between the maximum and minimum current density across the plane, Figure 4(a) shows how this variation changes with resistance (as in, with $R(y,z)=R_{0}$), Figure 4(b) shows how it changes with pouch cell length, and Figure 4(c) shows how it changes with current collector thickness. The figure identifies some near linear relationships between these cell parameters and the variation in current density, with explicit characterisations for these expressions obtainable from the analytical solution of Taheri, et al, 201432. As could be expected, the figure indicates that the variation of the current distribution across the pouch cell increases as: i) the cell resistance decreases, ii) the aspect ratio increases, and iii) the current collectors get thinner. Notably, Figure 4(a) indicates that as the cell ages, and so as its resistance increases, the variation in current distribution across the cell is likely to decrease, as more current is encouraged to flow along the current collectors than through the thickness of the cell.
@@
(b) Minimum, average, and maximum temperatures across the plane of the pouch cell for the experimental data and the model of (3) at 4C.
Figure 2: Comparison between the pouch cell temperature and voltage responses of the mathematical model of (3) and the experimental data of Lin, et al, 202220.
@@ 3 In-the-plane electrode grading
The spatial distributions seen in the response of the pouch cell, for example in the temperature data of Figure 3, presents an opportunity to apply electrode grading to create a more uniform electrochemical response. In this section, an analytic solution for the resistance distribution across the plane of the pouch cell, $R(y,z)$, to achieve a uniform current distribution is described.
@@ 3.1 Grading manufacturing
It is assumed that the electrode resistance $R(y,z)$ can be controlled locally in space. The reasons for focusing on the resistance distribution are: i) following Eqns. (3a) and (3b), the electrode resistance $R(y,z)$ directly influences the current distribution across the pouch cell (and it is variations in the current which are targeted to be smoothed out by the optimally graded electrode), ii) existing results on through-thickness electrode grading have demonstrated how this resistance grading can be achieved, e.g. by controlling the relative mass fraction of carbon black in LFP cathodes using spray printing techniques 15. It is envisaged that such spray techniques could be generalised to produce the optimised planar graded electrodes proposed in this study.
@@ 3.2 Resistance distribution for homogeneous current
Eqn. (3) implies that, under isothermal conditions of the open-circuit voltage, if a uniform current distribution is achieved at $t=0$, then $v_{1}$, $v_{2}$ and the SoC, will evolve uniformly in space as well. The implication is that if the current distribution is uniform at $t=0$, it will remain uniform during the simulation, following the model equations of (3a) and (3b), which may not hold exactly in practice as local current spikes might emerge caused by the variability in the local electrode composition. However, this can be used to simplify the graded electrode design optimisation problem. Specifically, it converts a dynamic optimisation problem into a static one. As in, it changes the optimisation from one of determining how the resistance should be distributed to achieve a uniform current during the whole charge to one that only has to consider the distribution at the start of the charge. A consequence of this problem formulation is that it allows an analytic solution to the graded electrode design optimisation to be derived (see Eqn (7)).
@@ 3.2.1 Problem formulation
With the above motivation, the desired uniform current distribution of Eqn. (5) is substituted into Eqns (3a) and (3b) to give, after simplifying,
$\Delta R+\frac{1}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}=0,\ \ \text{on}\ \ \Omega,$ (6a)
@@
$\frac{\partial R}{\partial z}\Big{|}_{z=L_{z}}$ $=-\frac{A_{\text{cell}}}{\sigma_{\scriptscriptstyle\text{ck}}A_{\scriptscriptstyle\text{tab},k}},\ \ \text{on}\ \ \partial\Omega_{\scriptscriptstyle\text{tab,k}},\,k\in\{n,p\}\ t>0,$ (6b)
@@ 3.2.2 Solution
$d_{2}$ $=-\frac{1}{2\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})},$ (7e)
@@
$h(y)=\begin{cases}\frac{L_{z}}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\text{ext}},~{}z=L_{z},\\ -\frac{A_{\text{cell}}}{\sigma_{\scriptscriptstyle\text{c1}}A_{\scriptscriptstyle\text{tab},1}}+\frac{L_{z}}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\scriptscriptstyle\text{tab,1}},~{}z=L_{z},\\ -\frac{A_{\text{cell}}}{\sigma_{\scriptscriptstyle\text{c2}}A_{\scriptscriptstyle\text{tab},2}}+\frac{L_{z}}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\scriptscriptstyle\text{tab,2}},~{}z=L_{z},\end{cases}$ (7g)
and the constant $d_{0}$ being a free variable setting the average value of the resistance.
@@
The solution of Eqn (7) shows that the resistance should roughly decrease quadratically away from the tabs in order to flatten the current distribution. The physical meaning of decreasing the resistance away from the tabs is to encourage current to flow down the length of the current collectors, instead of going directly through the electrodes as it enters one of the tabs and leaves the other. In the particular setup of Figure 1 and Table A2, where the cell has a relatively high aspect ratio, the variation of the resistance is much larger in the $z$-direction rather than the $y$-direction, but, following Eqn. (7), this may change for wider cell formats or for different tab configurations.
@@ 3.4 Results: Dynamic response of graded pouch cells
Figure 6 compares the optimised graded electrode cell following Eqn (7) and the uniform cell for a full 4C charge at $I_{\text{app}}$ = 80A. Figure 6 provides justification for the design choice of Section 3.1 to optimise the spatial resistance of the pouch cells to achieve a uniform initial current distribution in space, as Figure 6(a) indicates that this uniformity in the current was retained during the whole charge. By contrast, the current distribution in the uniform cell is more variable. Initially, large current gradients exist across the plane of the uniform cell, as shown in Figure 6(a), with the maximum current density 3925 A/m2 at the tabs and the minimum current 2138 A/m2 at the opposite end. The non-uniform current distribution then causes the voltage to vary, following Eqns. (3d)-(3f). Simultaneously, there is a negative feedback effect acting on the voltages from Eqns. (3a) and (3b) that smoothens this variation. The interactions between these two modes causes the spatial distribution of the current to evolve in both space and time. In particular, Figure 6(a) shows that the current does not change significantly during 100s $\leq t\leq$450s when the OCV curve is flat (since the smoothing effect of $U$ in Eqn. (3a) is not active). It is this smoothing effect which causes the region near the tabs to charge first, but then, when it is fully charged, the majority of the current then flows through the bottom half of the cell.
@@
Whilst Figure 6(a) indicates that there would be large differences between the current distributions of the uniform and graded cells, Figures 6(b) and 6(d) indicate that the impact of these differences on the variable internal to the current collectors would actually be quite limited. Specifically, Figure 6(b) compares the temperatures, Figure 6(c) the voltages, and Figure 6(d) the state-of-charges. Although there are some differences between the uniform and graded electrodes, it is relatively limited compared to the large variations seen in the current distributions. In particular, even though the profiles of the two voltage curves in Figure 6(c) are different, they reach the cut-off voltage of 3.85 V at roughly the same time (607s for the graded cell and 600s for the uniform one). As such, the capacitance improvement of the graded cell over the uniform one is only 1.2%. These limited gains in the capacitance are a result of the large thermal and electrical conductivities of the current collectors (see Table A2), which smooth out the variations in the electrodes responses across the plane. Similarly, for the current, Appendix Figure A1 compares the modelled 4C cycling results for the graded and uniform cells. The current density of the graded cell is close to $I(t)/A_{\text{cell}}$, while the uniform cell has large variations both in time and across the plane. The voltage of both cells is similar, while the maximum temperature is slightly reduced by $0.3\,^{\circ}\mathrm{C}$ for the graded cell.
@@
Figure 7: Current density distributions in the plane of the uniform pouch cell using the model of (3) during a constant current charge at 4C. Contour lines indicate current density values in A m-2 going through the electrode sandwich. The location of maximum current travels from the tabs to opposite them during the charge.
When only variables defined internally to the current collectors (such as temperature and voltage) are used to assess the value of graded pouch cells, their benefits may be lost. This observation agrees with the conclusions of Hosseinzadeh, et al, 201841 that porosity distribution through the electrode thickness has the potential to produce superior battery performance instead of when the porosity is varied along the electrode height. However, in this paper, it is proposed that to unlock the benefits of graded pouch cells, a performance metric that focuses on through-thickness effects should instead be considered. This observation motivates the analysis of the following section on the potential benefits of graded electrodes to mitigate lithium plating.
@@ 4 Results: Plating reduction with graded electrodes
The potential of graded electrodes to mitigate lithium plating in large format cells is explored in this section. As discussed in relation to Figures 6(b)-6(d) and in the results of Hosseinzadeh, et al, 201841, it was observed that the high conductivity of the current collectors limits the apparent value of electrode grading in-the-plane when only variables found “internally” to the current collectors (with internal variables being understood as those, such as voltage and temperature, that are smoothed out when flowing in the plane of the current collectors) are used to assess performance, as opposed to variables flowing through the electrodes (such as the current distributions of Figure 7 and 6(a)). However, as lithium plating is a localised degradation phenomenon dependent upon the charging current and electrochemical state of the electrode at a given point, it is shown in this section that plating may be effectively mitigated by in-the-plane electrode grading. The focus towards mitigating lithium plating is also motivated by experimental studies 24 which show significant spatial variations in plating across graphite anodes. Controlling the distribution of active carbon in the electrode microstructure to mitigate the spatial distribution in plating is proposed as a means to reduce the accelerated degradation of large format pouch cells in high-rate applications.
@@
Figure 8 compares the onset of lithium-ion plating for both the uniform and graded pouch cells using Eqn. (8). The purpose of this figure is to illustrate the spatial, temporal and C-rate rate dependence of plating, and how modelling can capture the heterogeneity of pouch cell degradation. Figure 8(a) shows the areal percentage of the pouch cell experiencing plating during charging at different C-rates. The uniform pouch cell sees a gradual increase in the area of the pouch cell experiencing plating when the C-rate is increased beyond 2.75C, with 100% of the pouch cell experiencing plating at 5.2C. By contrast, the graded pouch cell experiences no plating until 5.5C whereupon there is a sudden jump to 100% of its area that is plating. The differences between these two curves is due to variation in the current distribution of the uniform cell; this variation leads to higher current spikes and so a greater likelihood of plating. By contrast, the model predicts that the flatter current distribution of the graded pouch cell can delay the onset of plating until a critical point is reached, and, at that point, the flat distributions cause the whole plane to plate. The significance of this figure is that it suggests that grading pouch cells can increase the maximum charging C-rate for which plating can be avoided. In other words, grading in the plane could increase the operating envelope of pouch cells such that they could be fast charged without suffering from plating.

### v2

@@ 2.6 Model validation
Figure 2(b) shows that the model was able to capture the evolution of the maximum, minimum, and average values of the cell temperature across the plane of the pouch cell during the cycling. A main discrepancy between the results of the model and the experimental data is that the results for the model are much smoother. The increased variability of the data is to be expected, as the sensing noise and the inherent randomness in the microstructure introduced by manufacturing limitations are not captured by Eqn. (3). However, the model is still able to capture the general trends of the data, with the time constants and steady-state values of the first-order responses seen in the temperature dynamics being captured.
Figure 3 compares snapshots in the plane of the temperature distribution of the model and the experimental data 20, where the snapshots were taken at $t=$ 100s, 500s, 1000s, and 2500s. Even though the underlying model of Eqn. (3) is simple, strong agreement was made with the experimental thermal data from Lin et al., 202220. In particular, the formation of a thermal hot spot as well as the general shape of the temperature contours were captured by the model, although, again, the model’s diffusion dynamics resulted in smoother contours than those seen in the data. We note that Lin et al., 202220 presented a 3D model solved in COMSOL for the same data and a similar fit to that of Figure 3 was achieved. Therefore, even though the model of Eqn. (3) is arguably simpler than that of Lin et al., 202220 (see Section 2.4 for details), a similar level of model accuracy could be achieved.
Figure 4 illustrates how the validated model can be used to infer the impact of design changes in the pouch cell on the degree of spatial heterogeneity across the plane. Defining the variation in the current density as the difference between the maximum and minimum current density across the plane, Figure 4(a) shows how this variation changes with resistance (as in, with $R(y,z)=R_{0}$), Figure 4(b) shows how it changes with pouch cell length, and Figure 4(c) shows how it changes with current collector thickness. The figure identifies some near linear relationships between these cell parameters and the variation in current density (Figure 4 (b)), or its inverse (Figure 4 (a) and (c)), with explicit characterisations for these expressions obtainable from the analytical solution of Taheri et al., 201432. As could be expected, the figure indicates that the variation of the current distribution across the pouch cell increases as: i) the cell resistance decreases, ii) the aspect ratio increases, and iii) the current collectors get thinner. Notably, Figure 4(a) suggests that as the cell ages, and so as its resistance increases, the variation in current distribution across the cell is likely to decrease, as more current is encouraged to flow along the current collectors rather than through the thickness of the cell.
@@
(b) Comparison of the minimum, average, and maximum temperatures across the plane of the pouch cell between the experimental data and the model of (3) at 4C.
Figure 2: Comparison between the pouch cell temperature and voltage responses of the mathematical model of (3) and the experimental data of Lin et al., 202220.
@@ 3 In-the-plane electrode grading
The spatial distributions seen in the response of the pouch cell, for example in the temperature data of Figure 3, presents an opportunity to apply electrode grading to create a more uniform electrochemical response. In this section, an analytical solution for the resistance distribution across the plane of the pouch cell, $R(y,z)$, to achieve a uniform current distribution is described.
@@ 3.1 Grading manufacturing
It is assumed that the electrode resistance $R(y,z)$ can be controlled locally in space. The reasons for focusing on the resistance distribution are: i) following Eqns. (3a) and (3b), the electrode resistance $R(y,z)$ directly influences the current distribution across the pouch cell (and it is variations in the current which are targeted to be smoothed out by the optimally graded electrode), ii) existing results on through-thickness electrode grading have demonstrated how this resistance grading can be achieved, e.g. by controlling the relative mass fraction of carbon black in LFP cathodes using spray printing techniques 15. Achieving in-plane grading of electrodes using conventional techniques can be challenging. However, innovative methods such as spray printing show promise for creating a gradient across the thickness of electrodes. This technique can also effectively apply grading across the plane of the electrode. While spray printing has potential, it is currently limited to small-scale applications. In contrast, traditional methods like slurry casting can be creatively adapted to achieve partial in-plane grading8; 7. In future research, we will present experimental data on in-plane graded LFP electrodes fabricated using both spray printing and modified slurry casting techniques.
@@ 3.2 Resistance distribution for homogeneous current
Eqn. (3) implies that, under isothermal conditions of the open-circuit voltage, if a uniform current distribution is achieved at $t=0$, then $v_{1}$, $v_{2}$ and the SoC, will evolve uniformly in space as well. The implication is that if the current distribution is uniform at $t=0$, it will remain uniform during the simulation, following the model equations of (3a) and (3b), which may not hold exactly in practice as local current spikes might emerge caused by the variability in the local electrode composition. However, this can be used to simplify the graded electrode design optimisation problem, and smooth out current and temperature gradients across the pouch cells. Specifically, it converts a dynamic optimisation problem into a static one, in the sense that it changes the optimisation from one of determining how the resistance should be distributed to achieve a uniform current distribution during the whole charge to one that only has to consider the distribution at the start of the charge. A consequence of this problem formulation is that it allows an analytical solution to the graded electrode design optimisation to be derived (see Eqn (7)).
@@ 3.2.1 Problem formulation
The desired uniform current distribution of Eqn. (5) is substituted into Eqns (3a) and (3b) to give, after simplifying,
$\Delta R-\frac{1}{A_{\text{cell}}\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}=0,\ \ \text{on}\ \ \Omega,$ (6a)
@@
$\frac{\partial R}{\partial z}\Big{|}_{z=L_{z}}$ $=\frac{1}{\sigma_{\scriptscriptstyle\text{ck}}A_{\scriptscriptstyle\text{tab},k}},\ \ \text{on}\ \ \partial\Omega_{\scriptscriptstyle\text{tab,k}},\,k\in\{n,p\},$ (6b)
@@ 3.2.2 Solution
$d_{2}$ $=\frac{1}{2\;A_{\text{cell}}\;\sigma_{\text{eff}}\;(L_{\text{cn}}+L_{\text{cp}})},$ (7e)
@@
$h(y)=\begin{cases}-\frac{L_{z}}{A_{\text{cell}}\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\text{ext}},~{}z=L_{z},\\ \frac{1}{\sigma_{\scriptscriptstyle\text{c1}}A_{\scriptscriptstyle\text{tab},1}}-\frac{L_{z}}{A_{\text{cell}}\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\scriptscriptstyle\text{tab,1}},~{}z=L_{z},\\ \frac{1}{\sigma_{\scriptscriptstyle\text{c2}}A_{\scriptscriptstyle\text{tab},2}}-\frac{L_{z}}{A_{\text{cell}}\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})}&y\in\Omega_{\scriptscriptstyle\text{tab,2}},~{}z=L_{z},\end{cases}$ (7g)
and the constant $d_{0}$ being a free variable setting the average value of the resistance. This resistance distribution defines the graded electrodes of the following analysis. To evaluate Eqn. (7c), the infinite sum has to be truncated after a finite number of terms. It was found that a truncation of the order of 20 terms was sufficient to achieve a stable and accurate solution.
The solution of Eqn (7) shows that the resistance should roughly decrease away from the tabs in order to flatten the current distribution. The physical meaning of decreasing the resistance away from the tabs is to encourage current to flow down the length of the current collectors, instead of going directly through the electrodes as it enters one of the tabs and leaves the other. In the particular setup of Figure 1 and Table A2, where the cell has a relatively high aspect ratio, the variation of the resistance is much larger in the $z$-direction rather than the $y$-direction, but, following Eqn. (7), this may change for wider cell formats or for different tab configurations.
@@ 3.3 Results: Flat current distribution across pouch cells
Regarding whether the optimised resistance distributions of Figure 5 could be realised in practice, consider the maximum change in resistance that could be achieved by changing the local weight fraction of carbon black in the LFP cathodes. The characterisation of electronic conductivity of LFP cathodes in terms of the weight fraction of carbon black, $w_{\text{cb}}$, from Drummond et al., 15 is assumed. The resistance of the LFP cathode, $R_{\text{LFP}}$, is approximated15 by
$R_{\text{LFP}}=\frac{\ell}{n_{\text{layers}}\times A_{\text{cell}}\times\sigma_{\infty}\times\ {w_{\text{cb}}}^{b}}$ (8)
where $\sigma_{\infty}=4.01$ S m-1, $b=1.7$, and the electrode thickness is $\ell=100$ $\mu$m. Define $\overline{R}_{\text{LFP}}$ as the maximum resistance in Figure 5 (with a weight fraction of carbon black of $\overline{w}_{\text{cb}}$) and $\underline{R}_{\text{LFP}}$ as the minimum resistance (with a weight fraction of carbon black of $\underline{w}_{\text{cb}}$). Note that $\underline{w}_{\text{cb}}>\overline{w}_{\text{cb}}$ since carbon black increases the conductivity of the LFP electrode. The difference between the maximum and minimum resistances of the graded design is $\overline{R}_{\text{LFP}}-\underline{R}_{\text{LFP}}\approx 1.2\times 10^{-3}\,\Omega$. Therefore, the differences between the two weight fractions satisfies
$\frac{1}{{\overline{w}_{\text{cb}}}^{b}}-\frac{1}{{\underline{w}_{\text{cb}}}^{b}}=\frac{n_{\text{layers}}\times A_{\text{cell}}\times\sigma_{\infty}\times\left(\overline{R}_{\text{LFP}}-\underline{R}_{\text{LFP}}\right)}{\ell}.$ (9)
If the weight fraction of carbon black in the LFP cathode at the point with minimum resistance is $\underline{w}_{\text{cb}}=0.06$, then Eqn. 9 implies that the weight fraction at the point with maximum resistance should be $\overline{w}_{\text{cb}}=0.0471$. This is a modest difference and readily achievable in practice. Subtle changes in the electrode microstructure, for example in the distribution of carbon black in LFP cathodes, can therefore significantly influence how heterogeneities emerge across pouch cells.
@@ 3.4 Results: Dynamic response of graded pouch cells
Figure 6 compares the optimised graded electrode cell following Eqn (7) and the uniform cell for a full 4C charge at $I_{\text{app}}$ = 80A. Figure 6 provides justification for the design choice of Section 3.1, which optimises the spatial resistance of the pouch cells to achieve a uniform initial current distribution in space, as Figure 6(a) implies that this uniformity in the current was retained during the whole charge. By contrast, the current distribution in the uniform cell is more variable. Initially, large current gradients exist across the plane of the uniform cell, as shown in Figure 6(a), with the maximum current density 3925 A/m2 at the tabs and the minimum current 2138 A/m2 at the opposite end. The non-uniform current distribution then causes the voltage to vary, following Eqns. (3d)-(3f). Simultaneously, there is a negative feedback effect acting on the voltages from Eqns. (3a) and (3b) that smooths this variation. The interactions between these two modes causes the spatial distribution of the current to evolve in both space and time. In particular, Figure 6(a) shows that the current does not change significantly during 100s $\leq t\leq$450s when the OCV curve is flat (since the smoothing effect of $U$ in Eqn. (3a) is not active). It is this smoothing effect that causes the region near the tabs to charge first, but then, when it is fully charged, the majority of the current then flows through the bottom half of the cell.
@@
Whilst Figure 6(a) implies that there would be large differences between the current distributions of the uniform and graded cells, Figures 6(b) and 6(d) indicate that the impact of these differences on the variable internal to the current collectors would actually be quite limited. Specifically, Figure 6(b) compares the temperatures, Figure 6(c) the voltages, and Figure 6(d) the state-of-charges. Although there are some differences between the uniform and graded electrodes, it is relatively limited compared to the large variations seen in the current distributions. In particular, even though the profiles of the two voltage curves in Figure 6(c) are different, they reach the cut-off voltage of 3.85 V at approximately the same time (607s for the graded cell and 600s for the uniform one). As such, the capacitance improvement of the graded cell over the uniform one is only 1.2%. These limited gains in the capacitance are a result of the large thermal and electrical conductivities of the current collectors (see Table A2), which smooth out the variations in the electrodes responses across the plane. Similarly, for the current, Appendix Figure A1 compares the modelled 4C cycling results for the graded and uniform cells. The current density of the graded cell is close to $I(t)/A_{\text{cell}}$, while the uniform cell has large variations both in time and across the plane. By contrast, the maximum temperature is slightly reduced by $0.3\,^{\circ}\mathrm{C}$ for the graded cell.
@@
Figure 7: Current density distributions in the plane of the uniform pouch cell using the model of Eqn. (3) during a constant current charge at 4C. Contour lines indicate current density values in A m-2 going through the electrode sandwich. The location of maximum current travels from the tabs to opposite them during the charge.
When only variables defined internally to the current collectors (such as temperature and voltage) are used to assess the value of graded pouch cells, their benefits may be lost. This observation agrees with the conclusions of Hosseinzadeh et al., 201841 that porosity distribution through the electrode thickness has the potential to produce superior battery performance instead of when the porosity is varied along the electrode height. However, in this paper, it is proposed that to unlock the benefits of graded pouch cells, a performance metric that focuses on through-thickness effects should instead be considered. This observation motivates the analysis of the following section on the potential benefits of graded electrodes to mitigate lithium plating.
@@ 4 Results: Plating reduction with graded electrodes
The potential of graded electrodes to mitigate lithium plating in large format cells is explored in this section. As discussed in relation to Figures 6(b)-6(d) and in the results of Hosseinzadeh et al., 201841, it was observed that the high conductivity of the current collectors limits the apparent value of electrode grading in-the-plane when only variables found “internally” to the current collectors (with internal variables being understood as those, such as voltage and temperature, that are smoothed out when flowing in the plane of the current collectors) are used to assess performance, as opposed to variables flowing through the electrodes (such as the current distributions of Figure 7 and 6(a)). However, as lithium plating is a localised degradation phenomenon dependent upon the charging current and electrochemical state of the electrode at a given point, it is shown in this section that plating may be effectively mitigated by in-the-plane electrode grading. The focus towards mitigating lithium plating is also motivated by experimental studies 24 which show significant spatial variations in plating across graphite anodes. Controlling the distribution of active carbon in the electrode microstructure to mitigate the spatial distribution in plating is proposed as a means to reduce the accelerated degradation of large format pouch cells in high-rate applications.
@@
Figure 8 compares the onset of lithium-ion plating for both the uniform and graded pouch cells using Eqn. (10). The purpose of this figure is to illustrate the spatial, temporal and C-rate dependence of plating, and how modelling can capture the heterogeneity of pouch cell degradation. For this simulation, the temperature dependency of the onset of lithium plating was neglected as it was assumed that the temperature of the cell was approximately constant. This assumption could be relaxed using more complex plating models than Eqn. (10), such as those discussed in Reniers et al., 46, as the likelihood of plating is known to increase as the cell temperature decreases, especially for sub zero temperatureszinth2014lithium. Figure 8(a) shows the areal percentage of the pouch cell experiencing plating during charging at different C-rates. The uniform pouch cell sees a gradual increase in the area of the pouch cell experiencing plating when the C-rate is increased beyond 2.4C, with 100% of the pouch cell experiencing plating at 5.1C. By contrast, the graded pouch cell experiences no plating until 4.3C. The differences between these two curves is due to variation in the current distribution of the uniform cell; this variation leads to higher current spikes and so a greater likelihood of plating. By contrast, the model predicts that the flatter current distribution of the graded pouch cell can delay the onset of plating until a critical point is reached, and, at that point, the flat distributions cause the whole plane to plate. The predicted critical predicted C-rates for lithium plating are broadly consistent with the experimental literature; for example, Burns et al., burns2015situ detected plating in 220 mAh Li[Ni1/3Mn1/3Co1/3]O2/graphite 402035-size pouch cells at approximately 1C and 30 C. The significance of Figure 8(a) is that it suggests grading pouch cells could increase the maximum charging C-rate for which plating can be avoided. In other words, grading in the plane could increase the operating envelope of pouch cells such that they could be fast charged without suffering from plating.